# A Versatile Pathology Co-pilot via Reasoning Enhanced Multimodal Large Language Model

### v1

@@ 2 Results
Figure 7: Performance of MLLMs on WSI-Level VQA Tasks. a-b. Average performance of MLLMs based on accuracy (ACC) across 13 WSI-level VQA tasks. c-s. Model performance on specific tasks. * represents external validation datasets. Error bars represent 95% CI. The box limits represent the standard error. P-values are computed using a Wilcoxon signed-rank two-sided test. Additional results are shown in Extended Data Table 39 and Tables 40-43. t. An example of WSI-level VQA task along with the results generated by various MLLMs.
Unlike classification tasks where models select from a predefined set of labels, VQA tasks introduce a more flexible and clinically aligned framework for analyzing WSIs. In WSI-level VQA, models process pathology relevant questions about gigapixel WSIs and generate free-form answers or select from dynamic, context-dependent options. Accurately answering questions about WSIs requires the model to understand the global context, integrate information from different regions, apply domain knowledge, and handle ambiguity.
To evaluate the reasoning capabilities of SmartPath-R1 at the WSI level, we conducted a comparative analysis against existing MLLMs on 13 WSI-level VQA tasks (Figure 7). The results demonstrate that SmartPath-R1 achieves a significantly higher average ACC (0.695) across all tasks (Figures 7a-b) compared to other models (PathoR1: 0.505, Quilt-LLaVA: 0.350, LLaVA-Med: 0.309, Qwen2.5VL: 0.360), indicating superior overall performance. This suggests that SmartPath-R1 possesses a greater capacity for reasoning about complex histological patterns and integrating information across the entire WSI. Analysis of task-specific performance (Figures 7c-s) reveals that SmartPath-R1 exhibits particularly strong performance on questions related to histopathological changes, disease classification, grading, staging, treatment guidance (p < 0.001), suggesting its ability to reason about tumor-specific features in WSIs.
The comparative evaluation further presents SmartPath-R1’s critical advantage in reconciling diagnostic ambiguity, as exemplified in the squamous cell carcinoma (SCC) grading task (Figure 7t) where it correctly identifies the moderately differentiated (B) ground truth diagnosis. While other models produce binary interpretations (either well- or poorly-differentiated), SmartPath-R1 recognizes the nuanced histologic spectrum for moderate differentiation.
@@ 3 Discussion
The development of SmartPath-R1 represents a paradigm shift in computational pathology by addressing two critical limitations of existing multimodal large language models (MLLMs): (1) the lack of clinically meaningful reasoning capacity and (2) the artificial fragmentation of tasks across biological scales. Our results demonstrate that embedding sequential reasoning and multiscale integration into MLLMs significantly enhances diagnostic accuracy, interpretability, and clinical relevance.
@@
Multiscale Integration for Holistic Pathology Analysis. Clinical pathology requires simultaneous evaluation of subcellular details (ROI-level) and architectural patterns (WSI-level), which is absent in current MLLMs. SmartPath-R1 bridges this gap via dynamic mixture-of-experts mechanism, enabling adaptive switching of ROI-level morphometrics and WSI-level topological reasoning. This architecture eliminates the need for separate ROI-based and WSI models, streamlining workflows while improving diagnostic consistency.
@@
In conclusion, SmartPath-R1 establishes a new standard for AI in pathology by demonstrating that human-like diagnostic reasoning can be computationally modeled and optimized. By unifying multiscale analysis with reinforcement learning-driven decision-making, our framework transcends the limitations of current MLLMs, offering a scalable, interpretable, and clinically actionable solution. This work not only advances computational pathology but also provides a blueprint for developing next-generation diagnostic AI across medicine—where reasoning, not just recognition, determines success.
@@ 4.2 Details of Model Training
We design specialized reward functions for different critical pathological tasks: classification, VQA, detection and segmentation. As shown in Table 1, each task utilizes distinct pathological reasoning prompts : a system prompt $\mathsf{p^{sys}}$, and task-specific prompts ($\mathsf{p^{cls}}$,$\mathsf{p^{vqa}}$, $\mathsf{p^{det}}$, $\mathsf{p^{seg}}$). The model generates structured responses containing both reasoning process and final answer, which are parsed to extract task-specific outputs. All reward functions combine two key components:
@@
Segmentation Reward. Since MLLMs struggle to directly output pixel-level segmentation masks, we first provide spatial prompts in the form of bounding box coordinates, then leverage an off-the-shelf segmentation model MedSAM [34] to obtain the segmentation results. Specifically, the model outputs $\mathsf{r}=\operatorname{TP}(\mathbf{I},\mathsf{p^{sys}},\mathsf{p^{seg}})$ with parsed bounding boxes $\mathcal{P}=\{\mathbf{p_{i}}\}_{i=1}^{I}$. The reward combines:
@@ 4.3 Curated Datasets
This study leverages a diverse collection of histopathology datasets, categorized into ROI-level image-text pairs, ROI-level classification, ROI-level detection and segmentation, ROI-level VQA, WSI-level image-text pairs, WSI-level classification, and WSI-level VQA. Specifically, PathCap, PathInstruct, Quilt-1M, and SlideInstrction-Caption are used for supervised fine-tuning. The training sets from CCRCC, Chaoyang, CRC-MSI, ESCA, Pancancer-TIL, UniToPatho (ROI-level classification, see Extended Data Table 2 for classes), BCSS, CoCaHis, CoNIC2022, CRAG, DigestPath2019 Cell, DigestPath2019 Tissue, MIDOG Breast, MIDOG Neuroendocrine, Rings, SICAPv2, TNBC, WSS4LUAD (ROI-level detection and segmentation, see Extended Data Table 12 for classes), PathMMU, PathVQA (ROI-level VQA), TCGA (WSI-Level classification, see Extended Data Table 33 for classes), and SlideInstruction-VQA (WSI-Level VQA) are used for supervised fine-tuning and reinforcement fine-tuning. BreakHis, PanCancer-TCGA, GlaS, NuCLS, BRACS, CAMELYON, CPTAC-NSCLC, and BCNB are exclusively used for external validation to assess model generalizability.
@@
CCRCC (4 classes). The CCRCC dataset [35] comprises 52,723 annotated histopathology ROIs (300$\times$300 pixels) derived from WSIs of clear cell renal cell carcinoma (CCRCC) specimens. These ROIs are randomly sampled from two independent sources: the TCGA-KIRC repository and the Helsinki cohort. The dataset encompasses six distinct histological classes: malignant tumor regions (13,057 ROIs), normal renal parenchyma (8,652 ROIs), stromal tissue (5,460 ROIs), red blood cell accumulations (996 ROIs), non-informative background areas (16,026 ROIs), and heterogeneous tissue types including necrosis and artifacts (8,522 ROIs). For robust classification modeling, we focused exclusively on four biologically meaningful classes - cancer, normal tissue, stroma, and blood - excluding ambiguous and non-informative ROIs. The dataset is randomly partitioned into training (22,530 ROIs) and test (5,635 ROIs) sets while preserving class distributions. This curated subset facilitates precise evaluation of computational pathology algorithms while minimizing confounding factors from ambiguous labels. Comprehensive performance metrics are provided in Extended Data Table 4.
@@
Chaoyang (4 classes). The Chaoyang dataset [36] provides a comprehensive collection of histopathology ROIs for colorectal tissue analysis, comprising four clinically relevant classes: normal mucosa (1,816 ROIs), serrated lesions (1,163 ROIs), adenocarcinoma (2,244 ROIs), and adenoma (937 ROIs). All ROIs are standardized to 224$\times$224 pixels and divided into training (4,021 ROIs) and test (2,139 ROIs) sets using the official split to ensure reproducibility. Performance metrics are detailed in Extended Data Table 5.
@@
CRC-MSI (2 classes). The CRC-MSI dataset comprises 51,918 high-resolution histopathology ROIs (512$\times$512 pixels) derived from colorectal cancer specimens in the TCGA database [37]. Each ROI is annotated with patient-level Microsatellite Instability (MSI) status, categorized as either MSI-H (Microsatellite Instability-High) or NonMSIH (combining Microsatellite Instability-Low and Microsatellite Stable cases). Using the official dataset partition, we maintained rigorous separation between training (19,557 ROIs) and test (32,361 ROIs) sets to ensure unbiased evaluation. This binary classification framework enables development of AI models for MSI status prediction directly from histopathology images, a clinically important molecular characteristic in colorectal cancer. Performance metrics are reported in Extended Data Table 6.
@@
ESCA (11 classes). The ESCA dataset [38] comprises 367,229 histopathology ROIs (256$\times$256 pixels) extracted from 320 WSIs of esophageal adenocarcinoma and esophagogastric junction tumors, collected from four institutions: University Hospital Cologne (UKK, 22 slides), Landesklinikum Wiener Neustadt (WNS, 62 slides), TCGA (22 slides), and Charité - Universitätsmedizin Berlin (CHA, 214 slides). Each ROI is annotated with one of eleven histological classes: adventitia (71,131 ROIs), lamina propria mucosae (2,173 ROIs), muscularis mucosae (2,951 ROIs), muscularis propria (83,358 ROIs), regression tissue (56,490 ROIs), gastric mucosa (44,416 ROIs), esophageal mucosa (18,561 ROIs), submucosa (22,117 ROIs), submucosal glands (1,516 ROIs), tumor (63,863 ROIs), and ulceration (753 ROIs). For model development, we used the CHA dataset (178,187 ROIs) for training and combined UKK, WNS, and TCGA datasets (189,142 ROIs) for testing. All ROIs are resized to 224$\times$224 pixels for consistency. This multi-center dataset enables robust evaluation of AI models for esophageal carcinoma subtyping, with results reported in Extended Data Table 7.
@@
Pancancer-TIL (2 classes). We utilize the PanCancer-TIL dataset [39, 40] containing 304,097 histopathology ROIs (100$\times$100 pixels at 0.5 micrometers per pixel) for tumor-infiltrating lymphocyte classification. The dataset includes 54,910 TIL-positive ROIs (at least two TILs present in the image) and 249,190 TIL-negative ROIs, following the official train-val-test split (209,221:38,601:56,275 ROIs). All ROIs are resized to 256$\times$256 pixels for model input standardization. Performance metrics are reported in Extended Data Table 8.
@@
UniToPatho (6 classes). The UniToPatho dataset [41] provides 9,536 histopathology ROIs extracted from 292 WSIs of colorectal specimens, specifically annotated to support deep learning-based classification of colorectal polyps and adenoma grading. The dataset includes six clinically relevant classes: Normal tissue (950 ROIs), Hyperplastic Polyp (545 ROIs), Tubular Adenoma with High-Grade dysplasia (454 ROIs), Tubular Adenoma with Low-Grade dysplasia (3,618 ROIs), Tubulo-Villous Adenoma with High-Grade dysplasia (916 ROIs), and Tubulo-Villous Adenoma with Low-Grade dysplasia (2,186 ROIs). Following the official dataset partition, we used 6,270 ROIs for training and 2,399 ROIs for testing. This carefully curated collection enables development of AI models for precise polyp characterization, with experimental results detailed in Extended Data Table 9.
BreakHis* (2 classes). The BreakHis dataset [42] for breast cancer histopathological image classification is utilized in this study for external validation. The dataset comprises two primary classes: benign tumors and malignant tumors. All ROIs are captured at four distinct magnification levels (40$\times$, 100$\times$, 200$\times$, and 400$\times$). For consistency, images are resized to 224$\times$224 pixels. For external validation, only the test set is employed, consisting of 1,582 ROIs (20% of the total dataset), with stratification to preserve the original label distribution. Detailed experimental results are provided in Extended Data Table 10.
@@
PanCancer-TCGA* (32 classes). The PanCancer-TCGA dataset [43] is also employed for external validation in this study. It consists of 54,342 histopathology images (test set) extracted from WSIs, each resized to 256$\times$256 pixels and annotated across 32 distinct cancer types, including Head and Neck Squamous Cell Carcinoma, Bladder Urothelial Carcinoma, Uterine Carcinosarcoma, Colon Adenocarcinoma, Lymphoid Neoplasm Diffuse Large B-cell Lymphoma, Lung Squamous Cell Carcinoma, Brain Lower Grade Glioma, Esophageal Carcinoma, Pheochromocytoma and Paraganglioma, Sarcoma, Glioblastoma Multiforme, Adrenocortical Carcinoma, Uterine Corpus Endometrial Carcinoma, Prostate Adenocarcinoma, Breast Invasive Carcinoma, Stomach Adenocarcinoma, Pancreatic Adenocarcinoma, Skin Cutaneous Melanoma, Ovarian Serous Cystadenocarcinoma, Thymoma, Lung Adenocarcinoma, Kidney Renal Papillary Cell Carcinoma, Testicular Germ Cell Tumors, Kidney Renal Clear Cell Carcinoma, Rectum Adenocarcinoma, Cholangiocarcinoma, Cervical Squamous Cell Carcinoma and Endocervical Adenocarcinoma, Thyroid Carcinoma, Mesothelioma, Uveal Melanoma, Liver Hepatocellular Carcinoma, and Kidney Chromophobe. The test set is rigorously stratified to maintain class distribution, ensuring an unbiased assessment of model generalizability. Performance metrics, demonstrating superior results compared to baseline models, are detailed in Extended Data Table 11.
@@
BCSS. The Breast Cancer Semantic Segmentation (BCSS) dataset [44] is a comprehensive histopathology image dataset for breast cancer analysis derived from TCGA WSIs and annotated by expert pathologists through the Digital Slide Archive platform. It contains 9,192 annotated images uniformly resized to 1024$\times$1024 pixels, including 7,322 training images and 1,870 test images. The dataset provides detailed pixel-level annotations across 16 distinct tissue classes essential for pathological diagnosis: Mucoid material, Blood, Metaplasia NOS, Glandular secretions, Necrosis or debris, Plasma cells, Lymphatics, Blood vessel, Other immune, Dcis, Tumor, Stroma, Normal acinus or duct, Fat, Lymphocytic, and Angioinvasion.
@@
CoCaHis. CoCaHis comprises 82 hematoxylin-eosin stained frozen section images acquired intraoperatively from 19 patients with liver metastases of colon cancer. The dataset includes pixel-wise annotations created through a consensus process involving four board-certified pathologists, two pathology residents, and one final-year medical student, demonstrating substantial inter-rater agreement. Following the official data partitioning, we employed 58 images for training and 24 for testing, with all images standardized to 1024$\times$1024 pixels.
@@
CoNIC2022. The CoNIC2022 dataset [45] is utilized for nuclear segmentation in colorectal pathology collected across 16 institutions in three countries, with nuclei classified into six classes: epithelial cells, lymphocytes, plasma cells, neutrophils, eosinophils, and connective tissue cells (a composite class encompassing endothelial cells, fibroblasts, and muscle cells). Segmentation masks are generated using a semi-automated approach with subsequent manual refinement, yielding a combined dataset of 535,063 nuclei. While this hybrid annotation strategy enabled large-scale data curation, we acknowledge the potential for residual noise inherent in semi-automated methods, which we systematically evaluated through comparative pathologist review. All image ROIs are standardized to 1024$\times$1024 pixels and partitioned into 15,387 training images and 3,945 test images.
@@
CRAG. The CRAG dataset [46] contains 213 H&E colorectal adenocarcinoma WSIs for gland segmentation at 20x magnification from the University Hospitals Coventry and Warwickshire (UHCW) NHS Trust in Coventry, United Kingdom. We divide the slices into 1024$\times$1024 ROIs, resulting 1,429 and 321 images for training and testing, respectively.
@@
DigestPath2019 Cell. The DigestPath2019 Cell dataset [47] is clinically validated for signet ring cell carcinoma (SRCC) detection and segmentation, representing the first public resource specifically designed for this diagnostically challenging gastrointestinal cancer variant. The dataset is collected from 155 patients across gastric and intestinal mucosa, sourced from four leading Chinese medical institutions to ensure geographic and demographic diversity. Annotations underwent rigorous multi-tiered review by gastrointestinal pathologists, with senior experts adjudicating discordant cases to ensure label accuracy. All images are standardized to 1024$\times$1024 pixels to balance computational efficiency with diagnostic fidelity. The dataset is partitioned into 352 training and 85 testing images.
@@
DigestPath2019 Tissue. The DigestPath2019 Tissue dataset [47] consists of total 872 tissue sub-slices from 476 patients, which are extracted from both benign and malignant areas to cover as much variety of tissue appearance as possible. Two expert pathologists review the sub-images to ensure no uncertain tissues between benign and malignant. Then, the malignant images are manually annotated at pixel-level by four experienced pathologists, followed by expert pathologist examination. We divide the slices into 1024$\times$1024 images, resulting 10,725 and 2,666 images for training and testing, respectively.
MIDOG Breast. The MIDOG Breast dataset [48] comprises 150 WSI cases of human breast carcinoma evenly distributed across three digital scanning systems at 40$\times$ magnification obtained from the UMC Utrecht pathology archive. For computational analysis, these cases are processed into 3,794 annotated image ROIs, divided into a training set (3,032 ROIs) and an independent test set (762 ROIs). Each instance can be classified into one of two classes: breast cancer mitotic figure or breast cancer non-mitotic figures.
MIDOG Neuroendocrine. The MIDOG Neuroendocrine dataset [48] is a dataset for human pancreatic and gastrointestinal neuroendocrine tumor, characterized by their aggregated cellular morphology, obtained from the UMC Utrecht pathology archive. The dataset consisted of 55 WSI cases uniformly digitized using a Hamamatsu XR scanner at 40$\times$ magnification. We divide the slices into 1024$\times$1024 ROIs, resulting 1,448 and 354 images for training and testing, respectively.

### v2

@@ 2 Results
Figure 8: Performance of MLLMs on WSI-Level VQA Tasks. a. Average performance of MLLMs based on accuracy (ACC) across 13 WSI-level VQA tasks. b-r. Model performance on specific tasks. * represents external validation datasets. Error bars represent 95% CI. The box limits represent the standard error. P-values are computed using a Wilcoxon signed-rank two-sided test. Additional results are shown in Extended Data Table 39 and Tables 40-43. s. An example of WSI-level VQA task along with the results generated by various MLLMs.
Unlike classification tasks, where models select from a predefined set of labels, VQA tasks introduce a more flexible and clinically aligned framework for analyzing WSIs. In WSI-level VQA, models process pathology relevant questions about gigapixel WSIs and select from dynamic, context-dependent options. Accurately answering questions about WSIs requires the model to understand the global context, integrate information from different regions, apply domain knowledge, and handle ambiguity.
To evaluate the performance of SmartPath-R1 at the WSI level, we conduct a comparative analysis against existing MLLMs on 13 WSI-level VQA tasks (Figure 8). The results demonstrate that SmartPath-R1 achieves a significantly higher average ACC (0.695) across all tasks (Figures 8a) compared to other models (PathoR1: 0.505, Quilt-LLaVA: 0.350, LLaVA-Med: 0.309, Qwen2.5VL: 0.360), indicating superior overall performance. This suggests that SmartPath-R1 possesses a greater capacity for reasoning about complex histological patterns and integrating information across the entire WSI. Analysis of task-specific performance (Figures 8b-r) reveals that SmartPath-R1 exhibits particularly strong performance on questions related to histopathological changes, disease classification, grading, staging, and treatment guidance (p < 0.001), suggesting its ability to reason about tumor-specific features in WSIs.
The comparative evaluation further presents SmartPath-R1’s critical advantage in reconciling diagnostic ambiguity, as exemplified in the squamous cell carcinoma (SCC) grading task (Figure 8s), where it correctly identifies the moderately differentiated (B) ground truth diagnosis. While other models produce binary interpretations (either well- or poorly-differentiated), SmartPath-R1 recognizes the nuanced histologic spectrum for moderate differentiation.
@@ 3 Discussion
The development of SmartPath-R1 represents a paradigm shift in computational pathology by addressing two critical limitations of existing MLLMs: (1) the lack of clinically meaningful reasoning capacity and (2) the artificial fragmentation of tasks across biological scales. Our results demonstrate that embedding sequential reasoning and multiscale integration into MLLMs significantly enhances diagnostic accuracy, interpretability, and clinical relevance.
@@
Multiscale Integration for Holistic Pathology Analysis. Clinical pathology requires simultaneous evaluation of subcellular details (ROI-level) and architectural patterns (WSI-level), which is absent in current MLLMs. SmartPath-R1 bridges this gap via a dynamic mixture-of-experts mechanism, enabling adaptive switching of ROI-level morphometrics and WSI-level topological reasoning. This architecture eliminates the need for separate ROI-based and WSI models, streamlining workflows while improving diagnostic consistency.
@@
In conclusion, SmartPath-R1 establishes a new standard for AI in pathology by demonstrating that human-like diagnostic reasoning can be computationally modeled and optimized. By unifying multiscale analysis with reinforcement learning-driven decision-making, our framework transcends the limitations of current MLLMs, offering a scalable, interpretable, and clinically actionable solution. This work not only advances computational pathology but also provides a blueprint for developing next-generation diagnostic AI across medicine, where reasoning, not just recognition, determines success.
@@ 4.2 Details of Model Training
We design specialized reward functions for different critical pathological tasks: classification, VQA, detection, and segmentation. As shown in Table 1, each task utilizes distinct pathological reasoning prompts: a system prompt $\mathsf{p^{sys}}$, and task-specific prompts ($\mathsf{p^{cls}}$$\mathsf{p^{vqa}}$, $\mathsf{p^{det}}$, $\mathsf{p^{seg}}$). The model generates structured responses containing both the reasoning process and final answer, which are parsed to extract task-specific outputs. All reward functions combine two key components:
@@
Segmentation Reward. Since MLLMs struggle to directly output pixel-level segmentation masks, we first provide spatial prompts in the form of bounding box coordinates, then leverage an off-the-shelf segmentation model, MedSAM [29], to obtain the segmentation results. Specifically, the model outputs $\mathsf{r}=\operatorname{TP}(\mathbf{I},\mathsf{p^{sys}},\mathsf{p^{seg}})$ with parsed bounding boxes $\mathcal{P}=\{\mathbf{p_{i}}\}_{i=1}^{I}$. The reward combines:
@@ 4.3 Curated Datasets
This study leverages a diverse collection of histopathology datasets, categorized into ROI-level image-text pairs, ROI-level classification, ROI-level detection and segmentation, ROI-level VQA, WSI-level image-text pairs, WSI-level classification, and WSI-level VQA. Specifically, PathCap, PathInstruct, Quilt-1M, and SlideInstruction-Caption are used for supervised fine-tuning. The training sets from CCRCC, Chaoyang, CRC-MSI, ESCA, Pancancer-TIL, UniToPatho (ROI-level classification, see Extended Data Table 2 for classes), BCSS, CoCaHis, CoNIC2022, CRAG, DigestPath2019 Cell, DigestPath2019 Tissue, MIDOG Breast, MIDOG Neuroendocrine, Rings, SICAPv2, TNBC, WSS4LUAD (ROI-level detection and segmentation, see Extended Data Table 12 for classes), PathMMU, PathVQA (ROI-level VQA), TCGA (WSI-Level classification, see Extended Data Table 33 for classes), and SlideInstruction-VQA (WSI-Level VQA) are used for supervised fine-tuning and reinforcement fine-tuning. BreakHis, PanCancer-TCGA, GlaS, NuCLS, BRACS, CAMELYON, CPTAC-NSCLC, and BCNB are exclusively used for external validation to assess model generalizability. Importantly, to prevent potential data leakage, we meticulously screened and removed any samples from the original test sets that might have been inadvertently used during training (including potential overlaps across different groups of tasks). This precaution maintains the integrity of our validation process.
@@
CCRCC (4 classes). The CCRCC dataset [35] comprises 52,723 annotated histopathology ROIs ($\times300$ pixels) derived from WSIs of clear cell renal cell carcinoma (CCRCC) specimens. These ROIs are randomly sampled from two independent sources: the TCGA-KIRC repository and the Helsinki cohort. The dataset encompasses six distinct histological classes: malignant tumor regions (13,057 ROIs), normal renal parenchyma (8,652 ROIs), stromal tissue (5,460 ROIs), red blood cell accumulations (996 ROIs), non-informative background areas (16,026 ROIs), and heterogeneous tissue types including necrosis and artifacts (8,522 ROIs). For robust classification modeling, we focused exclusively on four biologically meaningful classes - cancer, normal tissue, stroma, and blood - excluding ambiguous and non-informative ROIs. The dataset is randomly partitioned into training (22,530 ROIs) and test (5,635 ROIs) sets while preserving class distributions. This curated subset facilitates precise evaluation of computational pathology algorithms while minimizing confounding factors from ambiguous labels. Comprehensive performance metrics are provided in Extended Data Table 4.
@@
Chaoyang (4 classes). The Chaoyang dataset [36] provides a comprehensive collection of histopathology ROIs for colorectal tissue analysis, comprising four clinically relevant classes: normal mucosa (1,816 ROIs), serrated lesions (1,163 ROIs), adenocarcinoma (2,244 ROIs), and adenoma (937 ROIs). All ROIs are standardized to $\times224$ pixels and divided into training (4,021 ROIs) and test (2,139 ROIs) sets using the official split to ensure reproducibility. Performance metrics are detailed in Extended Data Table 5.
@@
CRC-MSI (2 classes). The CRC-MSI dataset comprises 51,918 high-resolution histopathology ROIs ($\times512$ pixels) derived from colorectal cancer specimens in the TCGA database [37]. Each ROI is annotated with patient-level Microsatellite Instability (MSI) status, categorized as either MSI-H (Microsatellite Instability-High) or NonMSIH (combining Microsatellite Instability-Low and Microsatellite Stable cases). Using the official dataset partition, we maintained rigorous separation between training (19,557 ROIs) and test (32,361 ROIs) sets to ensure unbiased evaluation. This binary classification framework enables the development of AI models for MSI status prediction directly from histopathology images, a clinically important molecular characteristic in colorectal cancer. Performance metrics are reported in Extended Data Table 6.
@@
ESCA (11 classes). The ESCA dataset [38] comprises 367,229 histopathology ROIs ($\times256$ pixels) extracted from 320 WSIs of esophageal adenocarcinoma and esophagogastric junction tumors, collected from four institutions: University Hospital Cologne (UKK, 22 slides), Landesklinikum Wiener Neustadt (WNS, 62 slides), TCGA (22 slides), and Charité - Universitätsmedizin Berlin (CHA, 214 slides). Each ROI is annotated with one of eleven histological classes: adventitia (71,131 ROIs), lamina propria mucosae (2,173 ROIs), muscularis mucosae (2,951 ROIs), muscularis propria (83,358 ROIs), regression tissue (56,490 ROIs), gastric mucosa (44,416 ROIs), esophageal mucosa (18,561 ROIs), submucosa (22,117 ROIs), submucosal glands (1,516 ROIs), tumor (63,863 ROIs), and ulceration (753 ROIs). For model development, we used the CHA dataset (178,187 ROIs) for training and combined UKK, WNS, and TCGA datasets (189,142 ROIs) for testing. All ROIs are resized to $\times224$ pixels for consistency. This multi-center dataset enables robust evaluation of AI models for esophageal carcinoma subtyping, with results reported in Extended Data Table 7.
@@
Pancancer-TIL (2 classes). We utilize the PanCancer-TIL dataset [39, 40] containing 304,097 histopathology ROIs ($\times100$ pixels at 0.5 micrometers per pixel) for tumor-infiltrating lymphocyte classification. The dataset includes 54,910 TIL-positive ROIs (at least two TILs present in the image) and 249,190 TIL-negative ROIs, following the official train-val-test split (209,221:38,601:56,275 ROIs). All ROIs are resized to $\times256$ pixels for model input standardization. Performance metrics are reported in Extended Data Table 8.
@@
UniToPatho (6 classes). The UniToPatho dataset [41] provides 9,536 histopathology ROIs extracted from 292 WSIs of colorectal specimens, specifically annotated to support deep learning-based classification of colorectal polyps and adenoma grading. The dataset includes six clinically relevant classes: Normal tissue (950 ROIs), Hyperplastic Polyp (545 ROIs), Tubular Adenoma with High-Grade dysplasia (454 ROIs), Tubular Adenoma with Low-Grade dysplasia (3,618 ROIs), Tubulo-Villous Adenoma with High-Grade dysplasia (916 ROIs), and Tubulo-Villous Adenoma with Low-Grade dysplasia (2,186 ROIs). Following the official dataset partition, we used 6,270 ROIs for training and 2,399 ROIs for testing. This carefully curated collection enables the development of AI models for precise polyp characterization, with experimental results detailed in Extended Data Table 9.
[…]
@@
PanCancer-TCGA* (32 classes). The PanCancer-TCGA dataset [43] is also employed for external validation in this study. We utilize 2,000 histopathology images across 32 distinct cancer types, including Head and Neck Squamous Cell Carcinoma, Bladder Urothelial Carcinoma, Uterine Carcinosarcoma, Colon Adenocarcinoma, Lymphoid Neoplasm Diffuse Large B-cell Lymphoma, Lung Squamous Cell Carcinoma, Brain Lower Grade Glioma, Esophageal Carcinoma, Pheochromocytoma and Paraganglioma, Sarcoma, Glioblastoma Multiforme, Adrenocortical Carcinoma, Uterine Corpus Endometrial Carcinoma, Prostate Adenocarcinoma, Breast Invasive Carcinoma, Stomach Adenocarcinoma, Pancreatic Adenocarcinoma, Skin Cutaneous Melanoma, Ovarian Serous Cystadenocarcinoma, Thymoma, Lung Adenocarcinoma, Kidney Renal Papillary Cell Carcinoma, Testicular Germ Cell Tumors, Kidney Renal Clear Cell Carcinoma, Rectum Adenocarcinoma, Cholangiocarcinoma, Cervical Squamous Cell Carcinoma and Endocervical Adenocarcinoma, Thyroid Carcinoma, Mesothelioma, Uveal Melanoma, Liver Hepatocellular Carcinoma, and Kidney Chromophobe. The test set is rigorously stratified to maintain class distribution, ensuring an unbiased assessment of model generalizability. Performance metrics, demonstrating superior results compared to baseline models, are detailed in Extended Data Table 11.
@@
BCSS. The Breast Cancer Semantic Segmentation (BCSS) dataset [44] is a comprehensive histopathology image dataset for breast cancer analysis derived from TCGA WSIs and annotated by expert pathologists through the Digital Slide Archive platform. It contains 9,192 annotated images uniformly resized to $\times1024$ pixels, including 7,322 training images and 1,870 test images. The dataset provides detailed pixel-level annotations across 16 distinct tissue classes essential for pathological diagnosis: Mucoid material, Blood, Metaplasia NOS, Glandular secretions, Necrosis or debris, Plasma cells, Lymphatics, Blood vessel, Other immune, Dcis, Tumor, Stroma, Normal acinus or duct, Fat, Lymphocytic, and Angioinvasion.
@@
CoCaHis. CoCaHis comprises 82 hematoxylin-eosin stained frozen section images acquired intraoperatively from 19 patients with liver metastases of colon cancer. The dataset includes pixel-wise annotations created through a consensus process involving four board-certified pathologists, two pathology residents, and one final-year medical student, demonstrating substantial inter-rater agreement. Following the official data partitioning, we employed 58 images for training and 24 for testing, with all images standardized to $\times1024$ pixels.
@@
CoNIC2022. The CoNIC2022 dataset [45] is utilized for nuclear segmentation in colorectal pathology collected across 16 institutions in three countries, with nuclei classified into six classes: epithelial cells, lymphocytes, plasma cells, neutrophils, eosinophils, and connective tissue cells (a composite class encompassing endothelial cells, fibroblasts, and muscle cells). Segmentation masks are generated using a semi-automated approach with subsequent manual refinement, yielding a combined dataset of 535,063 nuclei. While this hybrid annotation strategy enabled large-scale data curation, we acknowledge the potential for residual noise inherent in semi-automated methods, which we systematically evaluated through comparative pathologist review. All image ROIs are standardized to $\times1024$ pixels and partitioned into 15,387 training images and 3,945 test images.
@@
CRAG. The CRAG dataset [46] contains 213 H&E colorectal adenocarcinoma WSIs for gland segmentation at 20x magnification from the University Hospitals Coventry and Warwickshire (UHCW) NHS Trust in Coventry, United Kingdom. We divide the slices into $\times1024$ ROIs, resulting in 1,429 and 321 images for training and testing, respectively.
@@
DigestPath2019 Cell. The DigestPath2019 Cell dataset [47] is clinically validated for signet ring cell carcinoma (SRCC) detection and segmentation, representing the first public resource specifically designed for this diagnostically challenging gastrointestinal cancer variant. The dataset is collected from 155 patients across gastric and intestinal mucosa, sourced from four leading Chinese medical institutions to ensure geographic and demographic diversity. Annotations underwent rigorous multi-tiered review by gastrointestinal pathologists, with senior experts adjudicating discordant cases to ensure label accuracy. All images are standardized to $\times1024$ pixels to balance computational efficiency with diagnostic fidelity. The dataset is partitioned into 352 training and 85 testing images.
@@
DigestPath2019 Tissue. The DigestPath2019 Tissue dataset [47] consists of total 872 tissue sub-slices from 476 patients, which are extracted from both benign and malignant areas to cover as much variety of tissue appearance as possible. Two expert pathologists review the sub-images to ensure no uncertain tissues between benign and malignant. Then, the malignant images are manually annotated at pixel-level by four experienced pathologists, followed by expert pathologist examination. We divide the slices into $\times1024$ images, resulting 10,725 and 2,666 images for training and testing, respectively.
[…]